OpenStudy (anonymous):

Two forces with magnitudes of 150 and 100 pounds act on an object at angles of 30° and 120° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (rottenywrm):

The resultant force on the object would be equal to the "dot product" of the two forces acting, which is Force(F) = 100 * 150 * cos(90) = 0. So there is not net force acting the object.

OpenStudy (anonymous):

where are you getting the cos(90) from?

OpenStudy (anonymous):

@rottenywrm

OpenStudy (anonymous):

@adrian_faz

OpenStudy (rottenywrm):

@Ziploxk "90" is the angle between the two forces. Resultant Force(F) = F1 * F1 * cos(Q) where Q is the smallest angle between the two forces and is also know as dot product of vectors.

OpenStudy (anonymous):

Okay but how was it found?

OpenStudy (rottenywrm):



OpenStudy (anonymous):

Okay can you walk me through the problem from start to finish. I have to be forgetting something.

OpenStudy (anonymous):

@adrian_faz

OpenStudy (mathmale):

Two forces are given; their directions are represented by the angles 30 and 120 degrees. What is the difference 120-30 degrees?